OpenStudy (anonymous):

subtract: (5x-6y+8)-(2x-3y+5)

OpenStudy (anonymous):

3x-3y+3 3(x-y+1)

Parth (parthkohli):

=5x - 6y + 8 - 2x + 3y - 5 =3x - 3y + 3 = 3(x - y + 1)